[figure: 9.7 cm]

ELECTRONIC DENSITY OF STATES OF YTTRIUM IRON GARNET

Ryszard J. Wojciechowski^{a},A. Lehmann-Szweykowska^{a}, Roman Micnas^{a},

and Gillian A. Gehring^{b}

aA. Mickiewicz University, Institute of Physics, ul. Umultowska 85, 61-614 Poznań,

Poland

bDepartment of Physics and Astronomy, University of Sheffield, Sheffield S3 7RH, UK

We have calculated the electronic structure of Fe^{3+}O_{4}^{2−} (tetrahedral) and Fe^{3+}O_{6}^{2−} (octahedral) clusters. Molecular-orbital point-symmetry cluster calculations have proven useful for providing insight into the real-space crystal unit cell consisting of 20 ironoxygen clusters. Discrete energy levels calculated with the point symmetry group S_{4} and S_{6}, respectively, are broadened due to charge-particle symmetry permitted inter-cluster motion. The transfer probability of spin up or down hole is energetically inequivalent. The on-site Coulomb repulsion on the iron ions is responsible for the spin dependance of the charge transport. The electronic DOS turns out to consist of narrow-bands.

[1] C. Srintiwarawong and G.A. Gehring, J. Phys.: Condens. Matter 13, 7987 (2001).

[2] A. Lehmann-Szweykowska, T. Lulek and M.M. Kaczmarek, J. Phys.: Condens. Matter 13, 3607 (2001).

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : oral

Corresponding author : Ryszard J. Wojciechowski

Address for correspondence : A. Mickiewicz University, Institute of Physics ul. Umultowska 85, 61-614 Poznań, Poland

Email address : wojrysz@main.amu.edu.pl